OpenStudy (anonymous):

Find the point on the terminal side of θ = -3 pi/4 that has an x coordinate of -1.

jimthompson5910 (jim_thompson5910):

Hint: The angle -3pi/4 radians is equivalent to (-3pi/4)*(180/pi) = -135 degrees which is coterminal to 360+(-135) = 225 degrees The reference angle is then 225 - 180 = 45 degrees since the angle is in Q III So this is what the drawing would look like 

jimthompson5910 (jim_thompson5910):

Your goal is to find the value of y of the point (-1, y)